# TAP4LLM: Table Provider on Sampling, Augmenting, and Packing Semi-structured Data for Large Language Model Reasoning

### 14 Dec 2023 v1

@@ 4.5. Comparison Results of Statistics Features
4.6. Trade-offs between token allocation
We use five table datasets to find the trade-off between token allocation for table sampling and table augmentation, as demonstrated in Figure 3. We find that:
(1) A balanced token distribution between the table and augmentation (around the 5:5 and 4:6 ratios, also known as a balanced T:A Ratio) generally yields the best performance for all five datasets. In indicates that properly controlling the token allocation can help LLMs achieve better performance.
Table 9. Ablation results on five table datasets using gpt-3.5-turbo model. Similar to Table 5, the lowest accuracy on each dataset is bold. The top-3 decreasing gap (delta) on each dataset are highlighted in red, with the lowest performance being underlined.
Components of TAP4LLM SQA FEVEROUS TabFact HybridQA ToTTo
Acc Delta Acc Delta Acc Delta Acc Delta BLEU-4 Delta
All 34.12% 0.00% 68.32% 0.00% 64.78% 0.00% 27.87% 0.00% 54.93% 0.00%
Figure 3. Token Allocation. T:A refers to the ratio of upper #token limitations of sampled table vs. augment info.
(2) Diminishing returns are observed when too many tokens are allocated to the augmentation information (as in the 3:7 ratio). This leads to a decrease in performance, suggesting that beyond a certain point, additional augmentation tokens may not be beneficial and could potentially detract from the main table content.
4.7. Ablation Study of TAP4LLM
As indicated in Table 9, we conduct an ablation study to assess the impact of various components on the performance of TAP4LLM. Each row represents the method performance without a certain component. We find that each component contributes to the model’s effectiveness. The study demonstrates that certain components, such as table sampling and table augmentation, are more crucial. Each dataset reacts differently to the removal of features, which highlights the necessity of a customized design when optimizing for particular datasets.
Figure 4. Comparative Analysis of Model Performance Across Four Datasets: FEVEROUS, HybridQA, TabFact, and SQA. The series of graphs illustrates the frequency distribution of token lengths alongside the LLM performance (%) for three distinct methods: only sampling, only augmentation, and the hybrid method. Each subplot corresponds to a different dataset, depicting how table token length impacts model accuracy for various data augmentation and sampling strategies
Figure 4. Comparative Analysis of Model Performance Across Four Datasets: FEVEROUS, HybridQA, TabFact, and SQA. The series of graphs illustrates the frequency distribution of token lengths alongside the LLM performance (%) for three distinct methods: only sampling, only augmentation, and the hybrid method. Each subplot corresponds to a different dataset, depicting how table token length impacts model accuracy for various data augmentation and sampling strategies
4.8. Large Table Analysis
Compared to the smaller-sized table, large table can grow to immense sizes, which make it more difficult to efficiently maintain, and reason over the tables. In designing TAP4LLM, performance optimization in this scenario is critical. We plot the distribution regarding the token numbers from the table across the five datasets in Figure 4, and also illustrate the impact of token numbers on LLM performance for three distinct methods: “only sampling”, “only augmentation”, and the “hybrid method”. Note that the “only sampling” method refers to using only the sampling method, without any augmentation methods. Specifically, we use query-based sampling with column grounding for this method. The “only augmentation” method refers to adding only the augmentation information to the prompt, without using any sampling method. In this method, all the augmentation information is simply combined. For the hybrid method, both table sampling and table augmentation methods are used to produce a unified framework. We can observe that:
(1) Data Concentration at Shorter Token Lengths: All datasets show that a majority of the data samples contain a smaller number of tokens, indicated by the frequency distribution that skews towards the left of the graph. This trend suggests that most of the text entries in these datasets are relatively brief.
(2) Performance with Only Augmentation: The method involving only augmentation seems to yield better accuracy at shorter token lengths. This could imply that augmentation techniques are more effective for smaller tables, which may be due to a more focused and less noisy dataset enabling the model to learn and generalize better from less complex input.
(3) Performance with Only Sampling: In contrast, the method that relies only on sampling appears to be more effective for longer token lengths, which correspond to larger tables. This may be because sampling allows the model to handle a greater diversity of data and to process on more informed, longer sequences, thus enhancing its ability to understand and generate correct answers from larger table bodies.
(4) Performance with the Hybrid Method: The hybrid method, which combines both augmentation and sampling techniques, shows a more consistent and stable accuracy across different token lengths. Notably, there’s a slight increase in accuracy for larger tables with longer token lengths. This suggests that the hybrid method benefits from the strengths of both augmentation and sampling, providing a balanced approach that ensures stable performance across varying text lengths and complexities.
5. Implementation Details
6. Related Work
Large Language Models Usage Paradigms. With the development of Generative Pre-trained Transformers (GPTs) (Radford et al., 2018, 2019; Brown et al., 2020; OpenAI, 2023, 2022) and other common LLMs (Chen et al., 2023; Zeng et al., 2022; Touvron et al., 2023; Zhang et al., 2022), many creative ways are proposed: End-to-End, Chain-of-Thought, and Semantic Parsing/Code Generation. Specifically, End-to-End (Chen, 2022; Hoffmann et al., 2022) aims to leverage LLMs to generate final answers directly, often done by providing a task description with a few examples for in-context learning. While convenient to use, this approach has the disadvantage of being uninterpretable and relying solely on the LLMs’ training data, which can lead to a lack of robustness and sensitivity to slight input variations. Chain-of-Thought  (Wei et al., 2022; Kojima et al., 2022b; Gong et al., 2020) emphasizes breaking down complex reasoning into a series of intermediate steps. However, this method is unreliable and uncontrollable due to the potential for generated reasoning steps to be misaligned with the intended logical progression, or for incorrect or inconsistent answers to be produced when multiple questions are asked consecutively. Semantic Parsing/Code Generation (Cheng et al., 2023; Ye et al., 2023; Gemmell and Dalton, 2023) leverages LLMs to convert natural language queries into executable code or structured representations. This approach enables more precise control over the generated output and facilitates better interpretability. Despite its advantages, this approach has several shortcomings, such as the limited coverage of programming language grammar and semantics. This restricts the model’s capacity to handle complex programming tasks and may necessitate the use of additional techniques to effectively process out-of-domain queries. Zhoujun Cheng et al. (Cheng et al., 2023) propose one way to bridge the out-of-domain queries through LLMs. However, how to inject knowledge still lacks a thorough exploration and the trigger functions are still naive for complex queries.
Large Language Models for Tabular Data. Following the line of LLMs in natural language processing, researchers have also explored large models for various modalities like vision (Gong et al., 2023; Kirillov et al., 2023) and speech (Huang et al., 2023). From a technical standpoint, their ability to generate human-like text has opened new vistas of possibilities for processing tabular data. Nevertheless, it is non-trivial to directly employ the vanilla ChatGPT model in the tabular area for two reasons: (i)-Global Table Understanding: the GPTs are known to suffer from the limited token length and thus, they can not read a whole large table, making them hard to understand the global tabular information. (ii)-Generalized to Tabular Domain: Second, their training processes are tailored for natural languages and thus, they are less generalizable when handling tabular data.
There have been several works (Hu et al., 2023; Zhong et al., 2017; Li et al., 2023a, b) developed to integrate natural language for tabular data analysis. NL2SQL (Nature language to SQL) (Hu et al., 2023; Zhong et al., 2017; Li et al., 2023a) is a long-standing research topic that converts natural language to SQL commands that manipulate the relational database. Recently, SheetCopilot (Li et al., 2023b) explored languages to VBA (Visual Basic for Applications, an embedded script language for Microsoft Excel) command to benefit from a rich set of spreadsheet software functionalities. Compared to previous works, TAP4LLM acts as a dedicated pre-processor designed specifically to improve the efficiency and effectiveness of LLMs for tabular reasoning tasks. We examine the considerable influence of pre-processing on the LLMs’ performance when dealing with tabular data. We particularly emphasize the potential of table sampling to extract essential information from the original table, table augmentation to incorporate existing symbolic models and world knowledge, and table packing to manage token allocation for ICL. Our experiment results show that TAP4LLM significantly outperforms the baselines for table reasoning tasks. Beyond its remarkable performance, our system also offers a systematic framework to spur further research into the interpretation of structured data in LLMs.
Table Augmentation. Table augmentation is a technique used to improve the generalization performance and robustness of machine learning models. To enhance the performance and capabilities of LLMs in various domains, various explorations have been done to augment their knowledge grounding. It involves incorporating structured knowledge (Sui et al., 2023; Xie et al., 2022), commonsense knowledge (Bian et al., 2023; Ogundare and Araya, 2023; Shen et al., 2023; Guo et al., 2023), and analytical knowledge (He et al., 2023; Jena et al., 2022) into the pre-training and inference processes. For example, (Jena et al., 2022) proposes to semi-automatically transform existing tabular data to create diverse/inventive natural language inference instances for better zero-shot performance. (He et al., 2023) proposes a multi-tasking Metadata model that leverages field distribution and knowledge graph information to accurately infer analysis metadata for tables, and then demonstrates its deployment in a data analysis product for intelligent features. We follow the definition of statistical features from (He et al., 2023). Each term with the corresponding definition is shown in Table 10. TAP4LLM crowd-source previous table augmentation methods and integrate them for empirical study. This helps us understand which type of knowledge is essential for table reasoning in Language Models (LLMs). We consider this unified knowledge scope to extend the knowledge and reasoning capabilities of LLMs, which enables LLMs to handle a more extensive range of complex tasks with improved accuracy and enhanced contextual understanding.
Table 10. Detailed definition of statistics features.
@@ 6. Related Work
7. Conclusion
We present TAP4LLM, a pre-processor designed to enhance the effectiveness of LLMs in tabular reasoning tasks. Technically, our method paves the way for interactive table reasoning as a natural language-driven framework, allowing for different components as plugins. Through three core components: table sampling, table augmentation, and table packing & serialization, we address several major challenges in comprehensive table understanding. We believe that TAP4LLM has the potential to revolutionize table reasoning by enhancing table modeling and exploratory data analysis (EDA) and enabling various domains such as finance, transportation, scientific research,etc. Our study will be beneficial for table-based research and serve as an auxiliary tool to help better utilize LLMs on tabular-based/structured data-based tasks.

### 17 Feb 2024 v2

@@ E.2 Comparison Results of Statistics Features
Table 8: Detailed definition of statistics features.
@@
Table 9: Comparative results of various types of statistical features. The experiment setting is the same as Section 2. The highest performance of each dataset is bold.
Statistics Features Type SQA FEVEROUS TabFact HybridQA ToTTo Spider
Progression features 29.20% 64.26% 60.45% 25.11% 49.53% 77.47%
String features 28.56% 63.13% 61.38% 24.83% 48.29% 73.56%
Number range features 29.13% 62.18% 59.03% 24.53% 49.68% 76.32%
Distribution features 30.28% 66.34% 62.18% 24.76% 49.34% 79.14%
Statistics features 31.22% 66.51% 62.33% 26.13% 50.57% 80.94%
E.3 TAP4LLM in Open-source model
Beyond conducting experiments on GPT models. we also evaluate the effectiveness of TAP4LLM on two most updated LLMs: Llama-2-70B and Mixtral-8x7B. According to Table 10, we first evaluated direct inference on open-source models and then apply TAP4LLM to each model. The result demonstrates that TAP4LLM increases models’ performance on all five datasets. The experiment gives us insights on the advantages of TAP4LLM among general LLMs. We will also conduct experiments on other table-related pre-trained models e.g.TaBERT Yin et al. (2020) in our future work.
Table 10: Comparison of TAP4LLM and baseline on Open-source models. We refer to "Baseline" as directly inferring each task using the model. For TAP4LLM, we apply semantic sampling for table sampling module and Statistics Feature/D/M+SF/self-prompting for table augmentation module.
Model Name Methods SQA FEVEROUS TabFact HybridQA ToTTo
Llama-2-70B Baseline 19.02% 65.33% 63.45% 17.21% 21.08%
TAP4LLM 22.14% 69.20% 66.32% 23.15% 30.00%
Mixtral-8x7B Baseline 21.25% 61.32% 57.21% 21.01% 34.25%
TAP4LLM 24.18% 63.29% 58.80% 25.44% 37.79%
E.4 TAP4LLM in Database Application
Dataset. We test TAP4LLM effectiveness on Spider Yu et al. (2018). Spider is a cross-domain Text-to-SQL dataset as shown in Table 5. Each instance contains a natural language question, a specific database containing tabular information, and one corresponding SQL query.
Metric. We evaluate TAP4LLM on the development split Spider-dev which contains 1034 instances over 200 databases. We use the Execution Accuracy, followed by the original paper Yu et al. (2018), to compare the execution output of the predicted SQL query with golden SQL query.
Experiment As shown in Table 11 and Table 12, the experiment result demonstrates that LLMs achieve an overall higher model performance through TAP4LLM. Specifically, the execution accuracy reaches the highest through semantic-based sampling and D/M + SF augmentation.
Table 11: Comparative results of the table sampling methods on Spider.
Sampling Type Table Sampling Methods Execution Accuracy
Rule-based Sampling Random Sampling 74.58%
Evenly Sampling 72.03%
Content Snapshot Yin et al. (2020) 78.93%
Embedding-based Sampling Centroid-based Sampling 77.43%
Semantic-based Sampling 80.27%
 w/ Column Grounding 81.03%
Hybrid Sampling 78.94%
LLM-based Sampling LLM-Decomposer Ye et al. (2023b) 78.34%
- No sampling (GPT-3.5) 72.15%
No sampling (GPT-3.5, truncated) 68.47%
Table 12: Comparative results of table augmentation methods on Spider. We use semantic-based sampling method without augmentation as the baseline for table augmentation.
Augmentation Aspect Execution Accuracy
Baseline 80.27%
D/M + SF 82.45%
Statistic Feature 80.94%
Term Explaination (LLM-based) 80.48%
Term Explaination (Heuristics-based) 80.33%
Appendix F Implementation Details
F.1 Motivation of our Framework
Table Sampling: One primary challenges for tabular reasoning is that the full content of a table could be very long and noisy to be include in the prompt. Most LLMs have a limited input context window size (e.g., 4k tokens) in which an overlong table cannot fit it. For long tables that satisfy the length constraint, it can still lead to unnecessary computations (of LLMs on long prompt) and quality regressions (generation interfered by noisy input) when placing irrelevant table content (w.r.t. the task or query) in the prompt.
Table Augmentation: Another challenge is what additional/external knowledge could help LLMs better understand a table? The raw content of a table may contain ambiguous information (e.g., abbreviations, domain-specific terms, column type, etc) that requires further interpretation and clarification. We are motivated to propose table augmentation for 1) enhanced contextual understanding: by supplementing tables with metadata and attributes, we can achieve a more profound grasp of the table’s intrinsic structure and semantics and further enrich the tabular data; 2) bridging external knowledge gasps: tables alone might not encompass all the required information to provide comprehensive answers to certain queries. By retrieving external knowledge from reliable sources, e.g., Wikipedia, we can aid the language models in understanding the broader context of the query, leading to more informed and nuanced responses.
Table Packing: The desire to maintain efficient reasoning without changing the LLMs architecture motivates us to consider how to encode the table into a prompt? While sampling and grounding compress the table content, augmentation expands the prompt by adding more information. With a given token budget, one needs to find the balance to allocate available tokens between table content and augmented knowledge.
F.2 Table Syncing
F.3 Table Cleansing
Table cleansing is an independent step in tabular data prepossessing, especially when dealing with hierarchical tables. In the context of fine-grained in-context learning, where pre-trained generated model has to discern and process intricate patterns and relationships within datasets. The importance of clean and standardized tables cannot be overstated for two reasons: (1) Dirty or unorganized tabular data can mislead the models and impair the model’s performance; (2) Cleansed tables ensure uniformity, making them easier to compare, merge, or use in subsequent operations. For example, imagine a financial analyst case aiming to forecast a company’s stock price based on historical data. The corresponding table contains daily stock prices, trading volumes, and various financial indicators. If there are any missing certain values for certain days, or duplicate entries due to system glitches. Such inconsistencies may dramatically affect the forecasting performance. For instance, it might suggest a non-trading day or a sudden drop in stock price. Specially, the formal definition of table cleansing is: Given a table $T$ consisting of rows $R_{T}$ and columns $C_{T}$, table cleansing transforms $T$ into $T^{\prime}$ such that: (a) Cell and column name completeness: For every cell $c_{i,j}$ in $T$ where $i\in R_{T}$ and $j\in C_{T}$, if $c_{i,j}$ has a missing or null value, it is filled using contextual information (i.e., use the corresponding entire column $C_{j}$ of cell $c_{i,j}$ as the context). We utilize a separate “CallLLM" system $g(\cdot)$ to call a pre-trained language model to synthesize the missing value. The processing can be formulated as $c_{i,j}=g(C_{j})$. This ensures that gaps in the data don’t lead to misleading interpretations or missed patterns. (b) Duplicate data points removal: For every pair of rows $r_{m}$, $r_{n}$ and pair of columns $c_{p}$, $c_{q}$ in $T$, if $r_{m}=r_{n}$ or $c_{p}=c_{q}$ respectively, one from the pair is removed to eliminate duplication. (c) Format consistency: For every cell $c_{i,j}$ in $T$, the value conforms to a specific format, unit, or pattern.